A positive test charge of 5.00 E-5 C is placed in an electric field. The force on it is 0.751 N. The magnitude of the electric field at the location of the test charge is


1
Expert's answer
2021-05-05T10:52:22-0400

We know that formula of force on charge is given by-

F= qE where q = charge , E=electric field

so "E=\dfrac{F}{q}=\dfrac{0.751}{5\times 10^{-5}}" = "15020\ N\ C^{-1}"

electric field at the location E ="15020\ N \ C^{-1}"